OpenStudy (steve816):

Easy quick question: Is 1^0 1 or undefined?

OpenStudy (skullpatrol):

1

OpenStudy (steve816):

Ok thanks, then 1 more quick question, is\[\log_{1}1 \]1 or undefined?

OpenStudy (skullpatrol):

Write it as y = log... and use the definition :-)

OpenStudy (steve816):

?? I'm not trying to make it into a function

OpenStudy (skullpatrol):

$$\huge 1^y = 1$$ for what values of y is that^ true?

OpenStudy (steve816):

Oh, all real numbers

OpenStudy (shawn):

you can not have base of 1.

OpenStudy (osprey):

my guess is that ANYTHING to the power 0 is 1. A sort of "catch all cop out clause" in maths ... ? That Rubik's cube looks "interesting" ...

OpenStudy (skullpatrol):

$$\huge y = log_b x $$ $$\huge b^y = x$$

zepdrix (zepdrix):

I think you.. or someone.. had asked this question the other day.. and I was curious and found a neat answer on stackexchange. If you try to apply change of base,\[\large\rm \log_11=\frac{\ln1}{\ln1}\]Which gives us division by zero, so it's just a way to show that this is a problem.

OpenStudy (shawn):

log_1(1) = 1 but this is just a special case. In general, the base can not be 1 because y = b^x is an exponential function. If b = 1, it's no longer an exponential function

OpenStudy (shawn):

the base is only defined to be between 0 and 1 exclusively, or strictly bigger than 1.

OpenStudy (steve816):

@zepdrix ah, that also makes sense.

OpenStudy (steve816):

And also, for logs, the bases can be negative and 0 correct?

OpenStudy (skullpatrol):

Good question! @steve816 :D

OpenStudy (steve816):

Because when I try to graph \[\large \log_{-2} x\] on desmos, I don't see anything.

OpenStudy (skullpatrol):

What definition are you using for logs?

OpenStudy (steve816):

What do you mean definition?

OpenStudy (skullpatrol):

What defines the logarithm of a number?

OpenStudy (steve816):

Just the standard definition\[\large \log_{a} x =y~~~~then~~~~~a^y=x\]

OpenStudy (skullpatrol):

So starting with $$a^y = x$$ are there any restrictions on a, x, or y?

OpenStudy (steve816):

a: can be all real numbers y: can be all real numbers x: x can't be 0

OpenStudy (skullpatrol):

If a and x are positive numbers ( a is not equal to 1), then $$log_a x = y$$ if and only if $$a^y = x$$

OpenStudy (steve816):

So... can the base a be a negative number?

OpenStudy (skullpatrol):

Not according to the standard definition.

OpenStudy (reemii):

Usually, when "0" or "empty set" are special cases of some calculation, you want to define it as the "default value" that make the other properties work fine (if possible). That's why * \(x^0 = 1\) because the property \(x^{a+0} = x^a x^0 = x^a\) is still satisfied. * \(\sum_{i\in \varnothing} x_i = 0\), so that \(\sum_{i\in A\cup \varnothing} x_i = \sum_{i\in A} x_i + \sum_{i\in \varnothing} x_i = \sum_{i\in A} x_i\) * \(\prod{i\in \varnothing} x_i = 1\), so that \(\prod_{i\in A\cup \varnothing} x_i = \prod_{i\in A} x_i \times \prod_{i\in \varnothing} x_i = \prod_{i\in A} x_i\) The "default" values are found like that.